OpenStudy (animallover8):

The time required to finish a test in normally distributed with a mean of 80 minutes and a standard deviation of 15 minutes. What is the probability that a student chosen at random will finish the test in more than 95 minutes?

OpenStudy (animallover8):

82% 2% 34% 16%

jimthompson5910 (jim_thompson5910):

Do you have a TI calculator?

OpenStudy (animallover8):

No, if I did I would be able to do many more questions.

jimthompson5910 (jim_thompson5910):

ok we'll use a table then

jimthompson5910 (jim_thompson5910):

first convert the raw score x = 95 to a z-score use the formula z = (x-mu)/sigma

jimthompson5910 (jim_thompson5910):

in this case, mu = 80 and sigma = 15

OpenStudy (animallover8):

What is the "raw score"?

jimthompson5910 (jim_thompson5910):

x = 95 is the raw score

jimthompson5910 (jim_thompson5910):

it's the score on the test the z-score is the transformed score to the standard normal distribution

jimthompson5910 (jim_thompson5910):

z = (x-mu)/sigma z = (95-80)/15 z = ???

OpenStudy (animallover8):

So z=...1?

jimthompson5910 (jim_thompson5910):

yes z = 1

jimthompson5910 (jim_thompson5910):

ie, z = 1.00

jimthompson5910 (jim_thompson5910):

Now use a table like this one https://www.stat.tamu.edu/~lzhou/stat302/standardnormaltable.pdf locate the row that starts with `1.0` and find the column that has `.00` at the top the intersection of this row and column has the number ______ (fill in the blank)

OpenStudy (animallover8):

.84134

OpenStudy (animallover8):

I don't understand what this number actually is.

jimthompson5910 (jim_thompson5910):

this number represents the probability of getting a z-score less than 1.00 in notation form, we say `P(Z < 1.00) = 0.84134` so there's approx a 84.134 % of getting a z-score less than 1.00 in other words, there is a 84.134 % chance of getting a test score less than 95

jimthompson5910 (jim_thompson5910):

subtract 0.84134 from 1 to get the probability of getting a score larger than z = 1 or x = 95

OpenStudy (animallover8):

0.269... ?

jimthompson5910 (jim_thompson5910):

compute 1 - 0.84134

OpenStudy (animallover8):

Wait so how do you get the probability from there?

jimthompson5910 (jim_thompson5910):

1 - 0.84134 = 0.15866

jimthompson5910 (jim_thompson5910):

0.15866 = 15.866% which rounds to 16%

jimthompson5910 (jim_thompson5910):

we have some normal curve 

OpenStudy (animallover8):

OOOOOOH!

jimthompson5910 (jim_thompson5910):